The length of a rectangle is 9 feet less than twice its width. The perimeter is 54 feet. What are the width and length of the rectangle?
Answer by stanbon(75887)

The length of a rectangle is 9 feet less than twice its width. The perimeter is 54 feet. What are the width and length of the rectangle?
-------
Let width = W
Then length = L = 2W-9
------------------------
Perimeter = 2(L+W)
54 = 2(2W-9+W)
27 = 3W-9
3W = 36
W = 12 ft (width)
L = 2W-9 = 15 ft (length)
=============================
Cheers,
Stan H.
===========